A triathlete bikes 15 miles at 20 miles per hour, runs 5 miles at 6 miles per hour and swims 1 mile at 4 miles per hour. How long does the triathlon take her to complete? Wouldn't the first answer be 45 minutes the 2nd part 50minutes and the third part 15 minutes? Why then is the answer 3 hours and 5 minutes?
1
Expert's answer
2011-11-25T08:14:36-0500
Biling time is
T1 = S1/V1 = 15/20 = 0.75, or 45 minutes;
running time is
T2 = S2/V2 = 5/6, or 50 minutes;
and swimming time is
T3 = S3/V3 = 1/4 = 0.25, or 15 minutes.
At last, total time is
T = T1 + T2 + T3 = 0.75 + 5/6 + 0.25 = 1 + 5/6, or one hour and 50 minutes.
As we see, the answer is first part 45 minutes the 2nd part 50minutes and the third part 15 minutes and the total time one hour and 50 minutes.